A simple random sample of 64 concert tickets was drawn from a normal population. The mean and standard deviation of the sample were $120 and $25, respectively.

Estimate the population mean with 90% confidence.

a.

LCL =115.95

UCL = 124.05

b.

LCL =114.85

UCL = 125.14

c.

LCL =115.99

UCL = 124.01

d.

LCL =114.78

UCL = 125.22
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = $120

sample standard deviation = s = $25

sample size = n = 64

Degrees of freedom = df = n - 1 = 64-1= 63

At 90% confidence level the t is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.1

\alpha / 2 = 0.1/ 2 = 0.05

t\alpha /2,df = t0.05,63 = 1.669

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.669 * (25 / \sqrt 64)

= 5.22

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

120 - 5.22 < \mu < 120 + 5.22

114.78 < \mu < 125.22

(114.78,125.22)

LCL = 114.78

UCL = 125.22

Option d is correct.